OpenStudy (amistre64):

I measure the circumference of a circle to be 25 inches which gives me a volume of 85 in^3. I measure the circumference of another circle to be 25 inches which gives me a volume of 75 in^3. Why?

OpenStudy (unklerhaukus):

Circumference is 2 x π x radius, i do not know how you are calculating the volume of a circle

OpenStudy (amistre64):

oh, but i never said the volume of the circle is being measured :)

OpenStudy (jamesj):

Because you're measuring a sphere.

OpenStudy (unklerhaukus):

i do not understand

OpenStudy (amistre64):

different sphere would be plausible if I were measuring different parts of them, so that could be a reason why

OpenStudy (amistre64):

but most likely when i measure a sphere i would be comparing the volume associated with the diameters; so it prolly aint comparing spheres

OpenStudy (amistre64):

unkle, how do you obtain a volume by measureing a circle?

OpenStudy (unklerhaukus):

volume of a sphere = 4/3 π r ^3 \[V = 4/3 \times π r ^{3}\]

OpenStudy (anonymous):

ur question is wrong circle doesn't have volume

OpenStudy (amistre64):

right, but that would imply the same size spheres .. how about: